# Rethinking Causal Discovery Through the Lens of Exchangeability

### v1

@@ Abstract
Causal discovery methods have traditionally been developed under two distinct regimes: independent and identically distributed (i.i.d.) and timeseries data, each governed by separate modelling assumptions. In this paper, we argue that the i.i.d. setting can and should be reframed in terms of exchangeability, a strictly more general symmetry principle. We present the implications of this reframing, alongside two core arguments: (1) a conceptual argument, based on extending the dependency of experimental causal inference on exchangeability to causal discovery; and (2) an empirical argument, showing that many existing i.i.d. causal‐discovery methods are predicated on exchangeability assumptions, and that the sole extensive widely-used real‐world “i.i.d.” benchmark (the Tübingen dataset) consists mainly of exchangeable (and not i.i.d.) examples. Building on this insight, we introduce a novel synthetic dataset [^1] All our code is open-source and available in tiago.brogueira@github.com that enforces only the exchangeability assumption, without imposing the stronger i.i.d. assumption. We show that our exchangeable synthetic dataset mirrors the statistical structure of the real‐world ”i.i.d.” dataset more closely than all other i.i.d. synthetic datasets. Furthermore, we demonstrate the predictive capability of this dataset by proposing a neural‐network–based causal‐discovery algorithm trained exclusively on our synthetic dataset, and which performs similarly to other state‐of‐the‐art i.i.d. methods on the real‐world benchmark.
@@ 1 Introduction
The aim of scientific research is often to find causal relationships between certain variables of interest. This stems either from a desire to intervene in the system under study, (and not merely be able to make accurate statements over the data we observe) or from a goal of uncovering causal mechanisms underlying the observations, seeking a deeper understanding of the underlying phenomena (Pearl, 2009). Traditionally, these relationships are uncovered by performing experiments (known in this context as interventions). However, this can often be impossible or impractical (Glymour et al., 2019). Furthermore, if not all intervened variables are controlled for, this can lead to erroneous results (e.g., Simpson’s paradox (Ameringer et al., 2009)). In these situations, the need arises to learn causal relationships from just observational data. This growing field, which uses a purely data-driven approach to learn causal structure (represented using a graph), is called causal discovery.
@@
Causal discovery methods are typically divided into two main categories: those for independent and identically distributed (i.i.d.) data and those for time-series data (Hasan et al., 2023), as the latter inherently violate the i.i.d. assumption underlying static causal models (Günther et al., 2023). This work will not address time-series data. [^2] From this point onwards, we will often refer simply to causal discovery, but it should be interpreted as referring to the family of causal discovery focused on i.i.d. data. An important result in this field is that, when working with i.i.d. purely observational data and without further assumptions, one can only identify the causal structure up to the so-called Markov equivalence class (MEC), which is defined as the set of causal graphs that satisfy the same conditional independence properties (Spirtes et al., 2001). In particular, in the case of a pair of dependent variables, $X$ and $Y$, both causal directions ($X\rightarrow Y$, $Y\rightarrow X$) belong to the same MEC, as there are no conditional independence properties to be satisfied. Furthermore, with i.i.d. observations, everything not explained by the causal mechanism is considered to be noise, often framed into the formalism as an exogenous variable (Zanga et al., 2022).
@@
In general, identifying one of the elements of the MEC requires interventions, which in this context have been formalized by Pearl (Pearl, 2009). Formally, interventions correspond to modifications of the base structural causal model by altering or replacing the causal mechanisms (i.e., the incoming functions) for one or more chosen variables. Each such intervention defines a new “environment” whose data are still drawn i.i.d. from the intervened model .
The concept of exchangeability was first used in the field of causal discovery by Guo et al. (Guo et al., 2024). The main motivation arose from the fact that, by considering the data to be exchangeable, it would allow more complex dependencies between the data, which could then be explored, even in the bivariate case.
In this paper, we argue that the advantage of considering the data to be exchangeable lies far beyond this original motivation. More specifically, it stems from the fact that it is a fairer real-world representation. In order to do so, we begin by laying out the problem and the work done so far in this field in Section 2. Then, in Section 3, we explore what are the exact implications of assuming exchangeability instead of i.i.d., while giving the relevant arguments why. Afterwards, we dive into our synthetic dataset, how it was generated, and how it compares to others in Section 4. Lastly, in Section 5, we introduce a neural network trained on this dataset, which serves both as a valid causal discovery method and a verification of the assumptions made throughout this work.
2 Problem Setting and Related Work
@@ 2.1 Exchangeability
Definition: An exchangeable sequence of random variables is a finite or infinite sequence $X_{1},X_{2},X_{3},\ldots$, such that for any finite permutation $\pi$ of the position indices $\{1,\ldots,N\}$, the joint distribution of the permuted sequences is the same as that of the original: An exchangeable sequence thus verifies:
@@
Additionally, a sequence $X_{1},\ldots,X_{N}$ is partially exchangeable if there exists a partition of the indices such that permutations within each partition are exchangeable.
In essence, exchangeability is a notion of symmetry (Guo et al., 2024); informally, it states that the order in which the variables appear does not matter. Naturally, all independent and identically distributed (i.i.d.) sequences are exchangeable since $P(X_{1},\ldots,X_{N})=\prod_{i=1}^{N}P(X_{i})$. However, the opposite is not necessarily true, since exchangeability doesn’t imply either independence or an identical distribution of the variables. To better understand why, consider the Pólya urn setup: a sequence of random draws of black or white balls from an opaque urn with replacement, whose proportion of black and white balls is unknown. In this scenario, each ball updates our belief over the probability distribution inside the urn. For example, if our first three draws are white balls, then it is more likely that our next draw is white and vice versa. Therefore, it breaks the independence property of i.i.d. data. However, this sequence is still exchangeable, since any finite sequence of draws from the urn provides exactly the same information regardless of the order, and therefore the different draws are indeed exchangeable. Additionally, it is also important to understand what it means for a sequence not to be exchangeable: it implies that the order of the variables matters. In other words, the observed sequence represents a time series.
The most important result regarding exchangeability is given by the famous de Finetti Theorem, stated next:
De Finetti Theorem (De Finetti, 1931):
@@
$P(X_{1},\ldots,X_{n})=\int_{\Theta}\prod_{i=1}^{n}P(X_{i}\mid\theta)\,d\mu(\theta).$ (2)
In Equation (2), $d\mu(\theta)$ can be replaced by $p(\theta)d\theta$, if $\mu$ is absolutely continuous with respect to the Lebesgue measure on $\Theta$ (which is the only case of interest in this paper, and thus assumed in the sequel). The theorem has been extended in several ways, including to non-binary variables.
This theorem states that any sequence of exchangeable variables can be seen as a mixture of i.i.d. sequences. Each i.i.d. sequence is given by conditioning the data on the parameter $\theta$ and the mixture’s proportions are consequently defined by $p(\theta)$. This result agrees with the intuition from the Pólya urn examples. The case of i.i.d. data can be seen as the special scenario when $p(\theta)=\delta(\theta_{0})$. In this case, the expression simply becomes $p(X_{1},\ldots,X_{n})=\prod_{i=1}^{n}p(X_{i}\mid\theta_{0})$, which is obviously i.i.d..
Lastly, it is important to mention that it is the exact existence of this parameter $\theta$ that allows the expression of learned information about the system. For example, in the original Pólya urn, one can reframe each draw’s dependence on the previous results by stating that they are informative over this parameter and its probability distribution. Naturally though, the solution for it is not unique under an infinite sequence of exchangeable data. Nevertheless, it is possible to infer valid values for both $p(x_{i}\mid\theta)$ and $p(\theta)$ from an infinite exchangeable sequence. This has been seen by Bernardo and Smith for a sequence of a Bernoulli variable (Bernardo and Smith, 2009).
In Bayesian Theory, this $\theta$ is interpreted as a latent variable. Note that $p(\theta)$, which appears naturally through the theory of exchangeability, is exactly equivalent to our prior belief on $\theta$ in Bayesian theory. Therefore, it is not by mathematical need or lack of knowledge that Bayesian inference requires a prior on $\theta$. In fact, it stems from $\theta$ being a real latent variable which indeed follows a given probability distribution (Fortini and Petrone, 2025). It demonstrates that it is the belief that the observations are exchangeable, not any metaphysical belief about the true model, that underpins the use of Bayesian modeling involving i.i.d. observations conditioned on some unknown latent variable.
2.2 Causal Discovery
Causal discovery is the problem of finding the graph that represents the causal relationships between a collection of variables of interest (Pearl, 2009). Usually, this graph is considered to be directed and acyclic. The absence of cycles implies it cannot capture the nature of systems with feedback. This assumption is taken in order to simplify the problem, however, in some cases, causal discovery methods may still work even if the acyclicity assumption does not hold (Glymour et al., 2019). Alternatively, some methods have also been proposed that relax the assumption of acyclicity (Richardson, 2013; Lacerda et al., 2012). Additionally, both the faithfulness assumption, which means that every conditional independence property in the joint probability distribution corresponds to a separation property in the graph, and the Markov property, which states that any variable should be independent of all other variables in the graph when conditioned on its parents, should hold (Pearl, 2009).
To answer counterfactual questions (”What would have happened if…?”), one requires a more complete description of the causal system than just the information contained in a causal graph. This description can be expressed in its most general form as a structural causal model (SCM). If we restrict an SCM to a specific set of parametrized functions $f(\cdot,\cdot;\eta)$, representing the causal relationships, we obtain a functional causal model (FCM), which is defined by its DAG and the functions relating its variables,
$X_{i}=f(PA(X_{i}),\epsilon_{i};\eta_{i}),$ (3)
where $PA(X_{i})$ is the set of parents of $X_{i}$ in the DAG and each $\epsilon_{i}$ is an exogenous variable (Zanga et al., 2022).
As mentioned in Section 2, in the bivariate scenario, it is impossible to distinguish between the two possible causal directions without interventions or additional assumptions. Therefore, different methods are based on different assumptions and argue why these should only hold in the true causal direction. Some try to exploit the independence between different variables, for example, independence between noise and cause (Peters et al., 2010) or independence between noise and function mechanism (Janzing et al., 2012). Another family of assumptions is based on Occam’s razor, restricting the set of accepted causal mechanisms and choosing the causal direction that has a better fit. For example, Blöbaum et al. restrict the functions to be linear, Goudet et al. opt for a neural network with one hidden layer, while Dhir et al. use Bayesian model selection with priors on the functions to be more flexible (Blöbaum et al., 2018; Goudet et al., 2018; Dhir et al., 2023). Lastly, some methods are based on analysing the complexity of the model in each direction, assuming the factorized model in the true direction has a lower Kolmogorov complexity; however, since Kolmogorov complexity is not computable, it is necessary to resort to different proxies of this measure (Marx and Vreeken, 2017; Tagasovska et al., 2020).
Regrettably, the causal discovery research area is critically lacking in extensive real‐world datasets [^3] Here, “extensive” denotes the presence of many distinct examples. Although numerous multivariate datasets exist, each comprises only a single causal graph, which precludes the use of evaluation metrics such as accuracy., with the Tübingen dataset remaining the only one widely accepted in the literature (Mooij et al., 2016). It contains a total of 108 causes effect pairs with known ground truth based on expert knowledge. These pairs were gathered from 37 different domains; some examples are the altitude and temperature of cities, the horsepower and fuel consumption of cars, and the age and height of different people. Given the lack of good real-world benchmarks, it is common to also test new methods on synthetic datasets. For the bivariate problem in particular, there exist 4 commonly used datasets (CE-Cha, CE-Net, CE-Gauss, and CE-Multi) Guyon et al. (2019). These datasets (with the exception of CE-Cha) were designed having certain assumptions in mind (such as additive or multiplicative noise) and thus are usually used to test whether certain methods can detect the true causal direction under those specific hypotheses and not to assess the more general performance of the methods.
Lastly, causal representation learning (CRL) is an emerging field that extends traditional causal discovery to high-dimensional data by aiming to recover latent causal factors from raw observations (Schölkopf et al., 2021). It assumes the observations are generated by a set of low-dimensional latent variables that follow an SCM and seeks to learn neural encoders and decoders that map between the observations and these abstract causal variables. By embedding a causal structure into a deep learning architecture, CRL enables the model to learn disentangled and modular representations of the underlying data-generating process. Such representations are often interpretable and robust, as independent causal factors tend to remain invariant under interventions or distribution shifts, allowing model components to be reused or fine-tuned for new tasks.

### v2

# Rethinking Bivariate Causal Discovery Through the Lens of Exchangeability

@@ Abstract
Causal discovery methods have traditionally been developed under two different modeling assumptions: independent and identically distributed (i.i.d.) data and time series data. In this paper, we focus on the i.i.d. setting, arguing that it should be reframed in terms of exchangeability, a strictly more general symmetry principle. For that goal, we propose an exchangeable hierarchical model that builds upon the recent Causal de Finetti theorem. Using this model, we show that both the uncertainty regarding the causal mechanism and the uncertainty in the distribution of latent variables are better captured under the broader assumption of exchangeability. In fact, we argue that this is most often the case with real data, as supported by an in-depth analysis of the Tübingen dataset. Exploiting this insight, we introduce a novel synthetic dataset [^1] All our code is open-source and available at https://github.com/tiagobrogueira/Thesis-Causal-Discovery-In-Exchangeable-Data/tree/main that mimics the generation process induced by the proposed exchangeable hierarchical model. We show that our exchangeable synthetic dataset mirrors the statistical and causal structure of the Tübingen dataset more closely than other i.i.d. synthetic datasets. Furthermore, we introduce SynthNN, a neural‐network–based causal‐discovery method trained exclusively on the proposed synthetic dataset. The fact that SynthNN performs competitively with other state‐of‐the‐art methods on the real‐world Tübingen dataset provides strong evidence for the realism of the underlying exchangeable generative model.
@@ 1 Introduction
The aim of scientific research is often to find causal relationships between variables of interest. This stems either from a desire to intervene in the system under study (and not merely be able to make accurate statements over the data we observe) or from a goal of uncovering causal mechanisms underlying the observations, seeking a deeper understanding of the underlying phenomena (Pearl, 2009). Traditionally, these relationships are uncovered by performing experiments (known in this context as interventions). However, this can often be impossible, impractical, or unethical (Glymour et al., 2019). Furthermore, if not all intervening variables are controlled for, this can lead to erroneous results (e.g., Simpson’s paradox (Ameringer et al., 2009)). In these situations, the need arises to learn causal relationships from observational data alone. This growing field, which uses a purely data-driven approach to learn causal structure (represented using a graph), is called causal discovery.
@@
Causal discovery methods are typically divided into two main categories: those assuming independent and identically distributed (i.i.d.) data and those designed for time series data (Hasan et al., 2023), which will not be addressed in this paper [^2] From this point onward, we will often refer simply to causal discovery, but it should be interpreted as referring to the family of causal discovery approaches for data traditionally assumed to be i.i.d. (Günther et al., 2023). An important result in this field is that, when working with purely observational i.i.d. data and no further assumptions, one can only identify the causal structure up to the so-called Markov equivalence class (MEC). This is defined as the set of causal graphs that satisfy the same conditional independence properties (Spirtes et al., 2001). In particular, in the case of a pair of dependent variables, $X$ and $Y$, both causal directions ($X\rightarrow Y$, $Y\rightarrow X$) belong to the same MEC, as there are no conditional independence properties to be satisfied. Furthermore, under the i.i.d. observations assumption, everything not explained by the causal mechanism is considered as noise, often formalized as exogenous variables (Zanga et al., 2022).
@@
In general, identifying one of the elements of the MEC requires interventions (Pearl, 2009). Formally, interventions correspond to modifications of the base structural causal model by altering or replacing the causal mechanisms (i.e., the incoming functions) for one or more chosen variables. Each such intervention defines a new “environment” whose data are still drawn i.i.d. from the intervened model.
Recently, the concept of exchangeability was brought into the field of causal discovery by Guo et al. (Guo et al., 2024). The main motivation of those authors is the fact that, by considering the data to be exchangeable, this allows more complex dependencies between the data points to be explored, even in the bivariate case. By treating bivariate data as an exchangeable sequence rather than strictly i.i.d., one can analyze relationships across multiple observations. This broader perspective unlocks cross-sample conditional independencies that break the observational symmetry, enabling the identification of the causal direction. Those authors propose a “Causal de Finetti” theorem, which demonstrates that exchangeable data containing specific cross-sample conditional independencies naturally represent independent causal mechanisms.
In this paper, we argue that the advantage of considering the data to be exchangeable goes beyond the original motivation of Guo et al. (Guo et al., 2024). Specifically, we show that it is also a more realistic model of real-world cause-effect pairs. Moreover, we enrich the exchangeable model of Guo et al. with a hierarchical structure, improving the flexibility of the exchangeability-based approach to bivariate causal discovery and of the Causal de Finetti theorem.
The rest of the paper is organized as follows. In Section 2, we begin by laying out the problem and reviewing previous work. In Section 3, we define and argue for the new hierarchical exchangeable model that better represents real-world causal mechanisms. Section 4 describes a new synthetic dataset generated in accordance with the proposed hierarchical model and compares it to other datasets. Finally, Section 5 introduces a neural network trained exclusively on this synthetic dataset, which serves as a new valid causal discovery method, and, most importantly, as further validation of the proposed model and its assumptions and rationale.
2 Causal Discovery, Exchangeability, and Their Intersection
2.1 Causal Discovery
Causal discovery is the problem of determining the causal relationships among a collection of dependent random variables of interest, say $X_{(1)},...,X_{(d)}$, usually described using a structural causal model (SCM) (Pearl, 2009). Typically, an SCM includes a directed acyclic graph (DAG) that specifies, for each $X_{(i)}$, what is the collection of other variables that are directly causally related to $X_{(i)}$, called the parents of $X_{(i)}$ and denoted $Pa(X_{(i)})$. Functional causal models (FCM) are a particular type of SCM where the dependency of each $X_{(i)}$ on its parents takes the form
$X_{(i)}=f_{i}(Pa(X_{(i)}),\epsilon_{i}),$ (1)
where $\epsilon_{1},...,\epsilon_{d}$ is a collection of so-called exogenous variables, which are considered to be mutually independent (Pearl, 2009; Peters et al., 2017; Zanga et al., 2022)
Causal discovery from purely observational data is an important, diverse, and large research field, which would be impossible to comprehensively review here (see (Zanga et al., 2022), for a recent survey and pointers to a vast literature). In this paper, we focus on what is arguably the quintessential causal discovery problem, called bivariate causal discovery or the cause-effect problem (Guyon et al., 2019; Mooij et al., 2016). Given two dependent variables $X$ and $Y$ (in practice, samples thereof), the famous Reichenbach common cause assumption states that ”there is no dependency without causation;” that is, either there is a third variable that is a common cause for $X$ and $Y$, called a confounder, $X$ causes $Y$, or $Y$ causes $X$. If there are reasons to assume causal sufficiency (i.e., absence of an unobserved confounder) and no feedback (Mooij et al., 2016), the problem becomes simply to select between the two possible causal directions: $X$ causes $Y$ or $Y$ causes $X$.
As mentioned in Section 2, in the bivariate case, it is impossible to distinguish between the two possible causal directions without interventions or additional assumptions. In the purely observation scenario, different methods are based on different assumptions, which typically take the form of some property that holds in the true causal direction but not in the reverse direction. Some methods exploit the independence between different variables, for example, independence between the noise exogenous variable and the cause (Peters et al., 2010) or independence between noise and function mechanism (Janzing et al., 2012). Another family of assumptions is based on Occam’s razor, restricting the set of accepted causal mechanisms and choosing the causal direction that has a better fit. For example, Blöbaum et al. (Blöbaum et al., 2018) restrict the functions to be linear, Goudet et al. (Goudet et al., 2018) opt for a neural network with one hidden layer, while Dhir et al. use Bayesian model selection with priors on the functions to be more flexible (Dhir et al., 2023). Lastly, some methods are based on analysing the complexity of the model in each direction, assuming the model in the true direction has a lower Kolmogorov complexity; however, since Kolmogorov complexity is not computable, it is necessary to resort to different approximations of this measure (Marx and Vreeken, 2017; Tagasovska et al., 2020).
Regrettably, the causal discovery research area is critically lacking in extensive real‐world datasets [^3] Here, “extensive” denotes the presence of many distinct examples. Although numerous multivariate datasets exist, each comprises only a single causal graph, which precludes the use of evaluation metrics such as accuracy., with the Tübingen dataset being the only one widely accepted in the literature (Mooij et al., 2016). It contains a total of 108 cause-effect pairs with known ground truth based on expert knowledge. These pairs were gathered from 37 different domains; some examples are the altitude and temperature of cities, the horsepower and fuel consumption of cars, and the age and height of different people. Given the lack of good real-world benchmarks, it is common to also test new methods on synthetic datasets. For the bivariate problem in particular, four commonly used datasets are CE-Cha, CE-Net, CE-Gauss, and CE-Multi Guyon et al. (2019). These datasets (except CE-Cha) were designed under certain assumptions (such as additive or multiplicative noise) and thus are usually used to test whether certain methods can detect the true causal direction under those assumptions, not to assess the more general performance of the methods.
@@ 2.2 Exchangeability
The concept of exchangeability, deeply tied to the foundational representation theorem by de Finetti, corresponds to a probabilistic notion of (permutation) symmetry (De Finetti, 1931). Intuitively, it establishes the conditions under which the order in which a sequence of observations is collected is irrelevant.
Definition: An exchangeable sequence of random variables is a finite or infinite sequence $X_{1},X_{2},X_{3},\ldots$, such that for any finite permutation $\pi:\{1,\ldots,N\}\rightarrow\{1,\ldots,N\}$ of the position indices, the joint distribution of the permuted sequences is the same as that of the original:
@@
Additionally, a sequence is said to be partially exchangeable if there exists a partition of the indices such that permutations within each partition are exchangeable.
Naturally, all independent and identically distributed (i.i.d.) sequences are exchangeable, but, of course, the converse is not true. Also, as the name implies, time series are obviously not exchangeable sequences since the order of the samples plays a central role in modeling this type of data. The following example aims to illustrate the concept of exchangeability.
Exchangeable urn sequence example:
Consider two urns (denoted as $A$ and $B$) with different ratios of black and white balls to the total number of balls (denoted $\theta_{A}$ and $\theta_{B}$). At the beginning of the experiment, a fair coin is flipped to select one of the urns, but which urn was chosen is not revealed to the observer. Subsequently, balls are randomly drawn with replacement from the selected urn. Let $X_{i}\in\{0,\,1\}$ denote the $i$-th sample, with 0 and 1 denoting white and black balls, respectively. The sequence thus generated is exchangeable; for example,
$P(X_{1}=1,X_{2}=0)=\theta_{A}(1-\theta_{A})\tfrac{1}{2}+\theta_{B}(1-\theta_{B})\tfrac{1}{2}=P(X_{1}=0,X_{2}=1).$
However, the samples are not independent; for example, since $P(X_{t}=1)=\tfrac{1}{2}\theta_{A}+\tfrac{1}{2}\theta_{B}$, for any $t$,
$P(X_{1}=1)\,P(X_{2}=0)=(\tfrac{1}{2}\theta_{A}+\tfrac{1}{2}\theta_{B})(1-\tfrac{1}{2}\theta_{A}-\tfrac{1}{2}\theta_{B})\neq P(X_{1}=1,X_{2}=0),$
thus the sequence is not i.i.d. (except, of course, in the trivial case where $\theta_{A}=\theta_{B}$).
The most important result regarding exchangeability is given by the famous de Finetti representation theorem (De Finetti, 1931), stated next.
De Finetti theorem:
@@
$P(X_{1},\ldots,X_{N})=\int_{\Theta}\prod_{n=1}^{N}P(X_{n}|\theta)\,d\mu(\theta).$ (3)
In Equation (3), $d\mu(\theta)$ can be replaced by $p(\theta)\,d\theta$, if $\mu$ is absolutely continuous with respect to the Lebesgue measure on $\Theta$, i.e., if there is a probability density function $p(\theta))$. In the case where $\Theta$ is a discrete set, Equation (3) takes the form
$P(X_{1},\ldots,X_{N})=\sum_{\theta\in\Theta}\prod_{n=1}^{N}P(X_{n}|\theta)\,p(\theta),$ (4)
where, in this case, $p(\theta)$ is a probability mass function. The theorem has been extended in several ways, including to non-binary variables (Bernardo and Smith, 2009). In essence, this theorem states that any sequence of exchangeable variables can be seen as a mixture of i.i.d. sequences, each of which is conditioned on the parameter $\theta$.
Exchangeable urn sequence example (continued):
The example above illustrates the de Finetti theorem with $\Theta=\{\theta_{A},\,\theta_{B}\}\subset[0,1]$, $p(\theta_{A})=p(\theta_{B})=1/2$, and $P(X_{n}|\theta)=\theta^{X_{n}}\,(1-\theta)^{1-X_{n}}$.
The existence of the latent (unobserved) parameter $\theta$ mediates the dependence between different variables in an exchangeable sequence. In the urn experiment, for example, each draw provides information about this underlying parameter, which, in turn, informs our expectations about future draws. In Bayesian theory, $\theta$ is treated as a latent variable, and the distribution $p_{\Theta}(\theta)$, which emerges naturally from de Finetti’s theorem, acts as the Bayesian prior (Bernardo and Smith, 2009). Thus, the requirement for a prior in Bayesian inference does not stem from mathematical convenience or a lack of knowledge, but from the fundamental modeling of $\theta$ as a latent variable (Fortini and Petrone, 2025). Ultimately, this demonstrates that it is our judgment of exchangeability, not any metaphysical belief about a “true” model, that underpins the standard use of Bayesian modeling involving i.i.d. observations conditioned on an unknown latent variable.
2.3 Hierachical Models
In this subsection, we enrich the class of models described above with additional latent variables, yielding so-called hierarchical models (Fortini and Petrone, 2025). The rationale for the introduction of these hierarchical models is as follows: sometimes, a complicated distribution $P(X|\theta)$ can be conveniently represented as the marginal with respect to some latent variable $\eta$, $P(X|\theta)=\int P(X|\eta,\theta)p(\eta|\theta)\,d\eta$; mixture models are a classical example.
The complete family of models is given by the following set of instances of the de Finetti representation, which we will comment on in detail below:
$\text{Model 1:}\;\;P(X_{1},\ldots,X_{N}|\theta)$ $=\prod_{n=1}^{N}P(X_{i}|\theta),$ (5)
$\text{Model 2:}\;\;P(X_{1},\ldots,X_{N})$ $=\int_{\Theta}\prod_{n=1}^{N}P(X_{n}|\theta)\,p(\theta)\,d\theta,$ (6)
$\text{Model 3:}\;\;P(X_{1},\ldots,X_{N}|\theta)$ $=\prod_{n=1}^{N}\int_{H}P(X_{n}|\eta_{n})\,p(\eta_{n}|\theta)d\eta_{n},$ (7)
$\text{Model 4:}\;\;P(X_{1},\ldots,X_{N})$ $=\int_{\Theta}\left(\prod_{n=1}^{N}\int_{H}P(X_{n}|\eta_{n})\,p(\eta_{n}|\theta)d\eta_{n}\,\right)p(\theta)\,d\theta,$ (8)
$\text{Model 5:}\;\;P(X_{1},\ldots,X_{N}|\theta)$ $=\prod_{n=1}^{N}\int_{H}P(X_{n}|\eta_{n},\theta)\,p(\eta_{n}|\theta)d\eta_{n},$ (9)
$\text{Model 6:}\;\;P(X_{1},\ldots,X_{N})$ $=\int_{\Theta}\left(\prod_{n=1}^{N}\int_{H}P(X_{n}|\eta_{n},\theta)\,p(\eta_{n}|\theta)d\eta_{n}\,\right)p(\theta)\,d\theta.$ (10)
For simplicity of notation, we are assuming that all random variables have densities. In the case of discrete random variables, the corresponding integrals are replaced by sums. In Figure 1, we use probabilistic graphical models (Koller and Friedman, 2009) to describe the structure of these hierarchical models. In this figure, the plate (rectangle with subscript $N$) means that there are $N$ conditionally i.i.d. copies of the variables inside, given the variable outside. Gray nodes correspond to latent/unobserved variables, while white nodes correspond to observed/known variables.
In models 1, 3, and 5, parameter $\theta$ is known/observed, thus we write $P(X_{1},\ldots,X_{N}|\theta)$. These three models describe i.i.d. (thus exchangeable) sequences, with increasingly complex hierarchical structures. Model 1 corresponds to a simple i.i.d. sequence, where each $X_{n}\sim P(X|\theta)$. In model 3, we have $X_{1},...,X_{n}$ conditionally independent given $\eta_{1},...,\eta_{N}$, with $X_{n}\sim P(X|\eta_{n})$ and, in turn, the $\eta_{n}\sim p(\eta|\theta)$ conditionally i.i.d., given $\theta$. Consequently, $P(X_{n}|\theta)$ is given by the inner integral in Equation (7). Finally, model 5 is an generalization of model 3 where $X_{n}\sim P(X|\eta_{n},\theta)$, i.e., $X_{n}$ is not conditionally independent of $\theta,$ given $\eta_{n}$.
Models 2, 4, and 6 correspond to exchangeable (but not i.i.d.) sequences, as the outer variable $\theta$ in the hierarchy is not observed, thus it is marginalized out in Equations (6), (8), and (10). Model 2 is simply the one addressed in the previous subsection, as Equation (6) coincides with Equation (3). Models 4 and 6 include an additional collection of random variables, with a structure similar to models 3 and 5, respectively.
[…]
Figure 1: Graphical models representing the probabilistic models 1 to 6, defined in Equations (5) – (10). The plate represents NN i.i.d. copies of the random variable inside, conditioned on the one outside. The variables in white are observed, those in gray are not.
The set of graphical models presented in Figure 1 illustrates that it is the ignorance versus knowledge about the outer variable $\theta$ (not its existence) that makes the sequence exchangeable but not independent (models 2, 4, and 6) versus i.i.d. (models 1, 3, 5). Furthermore, the presence (models 3, 4, 5, and 6) or absence (models 1 and 2) of the latent variables $\eta_{1},...,\eta_{N}$ inside the plate has no effect of the independence or not of $X_{1},...,X_{n}$.
We conclude this subsection with another urn example illustrating the hierarchical exchangeable model.
Hierarchical exchangeable urn sequence example:
Consider two same two urns ($A$ and $B$) with different ratios of black and white balls to the total number of balls (denoted $\eta_{(A)}$ and $\eta_{(B)}$). Now, in each draw (rather than only once in the beginning of the process), a coin is flipped to select one of the urns (although the choice is not observed), and a ball is sampled from the selected urn, with replacement. This coin is one of two different coins, coin 1 and coin 2, selected before the sampling process begins by flipping a third, fair coin. Coin 1 selects urn A with probability $\theta_{1}$ (thus urn $B$ with probability $1-\theta_{1}$) and coin 2 selects urn A with probability $\theta_{2}$ (thus urn $B$ with probability $1-\theta_{2}$). Which coin is being used is not observed. This corresponds to a mixture of two i.i.d. processes, such that each sample of this process is itself also a mixture of two distributions (the two urns).
This process follows model 4 defined by Equation (8), with the following definitions: $X_{n}\in\{0,\,1\}$, corresponding to black and white balls, respectively, and $P(X_{n}|\eta_{n})=\eta_{n}^{X_{n}}(1-\eta_{n})^{(1-X_{n})}$, for $\eta_{n}\in\{\eta_{(A)},\theta_{(B)}\}$. $P(\eta_{(A)}|\theta_{1})=\theta_{1}$ and $P(\eta_{(A)}|\theta_{2})=\theta_{2}$. Finally, $p(\theta_{1})=p(\theta_{2})=1/2$.
Note that it is not the mere existence of latent variables that makes the problem exchangeable. If $\theta$ and $\psi$ can only take one value, the outer integrals can be dropped, and the sequence is i.i.d. by definition. In the example above, it would be equivalent to having a single coin with known probabilities. Simultaneously, it is not the existence of both urns, or the fact that the experimental procedure alternates between urns, that makes this sequence exchangeable. It is the fact that the probability of sampling from each urn is unknown (initial choice of coin) that makes it so.